Shift report 20011111 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Nov 11 2001 - 07:57:39 EST

  • Next message: Flemming Videbaek: "Report from sceduling physicist & STAR problems"

    ------------------------------------------------------------
    Date: Sun Nov 11 07:57:38 2001, Shift: 00:00-08:00
    Supervisor: TST, with: 
    Log book#10 pages 125 - 126
    
    Shift Summary:
    ------------------------------------------------------------
    1:  Store from previous shift continued until 03.45, with
    falling rates (voted with Phenix to dump.)  Attempts to
    inject towards end of shift.
    
    
    2:  Run    Trig6  
       5757     136k   MRS@90,4A, FFS@30,BFS@6,1/15B
       5758            DVM pedestal run
       5759            TPC pedestal run
       5760            Pedestal run
    
    3:  No new detector problems.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Nov 11 2001 - 07:58:15 EST